# When Worse is Better: Navigating the compression-generation tradeoff in visual tokenization

### v1

@@ Appendix C More Detailed Scaling Plots
Figure 13 expands on Figure 7, showing full model size scaling details and best fit scaling laws. We see the same trends discussed in the original plot, where 1k performs better in low regimes and 131k performs better in high compute regimes.
Figure 14 expands on Figure 4, showing scaling across three different tokens per image amounts on all of our model scales. We see the same trend but in more detail, where 576 performs poorly in low compute regimes (and with small models) but outperforms 400 and 256 tokens per image at high compute regimes. Note that 256 tokens per image is compute optimal up until the 2.5 gFID barrier is reached,at which point the tokenizer reconstruction quality becomes a bottleneck.
@@
Figure 12 provides a different perspective on the comparison between Figures 3 and 8. We do this by comparing stage 2 model scales directly by gFID performance by train FLOPs. Note that like the main figure, each point is an independent training run. We see that CRT tokens improve performance at all scales both for training efficiency and final performance. This difference is especially pronounced at the smaller compute and parameter scales.
Appendix D Qualitative Examples
@@
Figures 15-19 show non-cherrypicked generation examples from our best model CRTopt-AR-775M. They demonstrate a high degree of diversity and image quality across a wide array of classes.

### v2

@@ Appendix F More Detailed Scaling Plots
Table 5: Text-to-image results on BLIP-3o. Re-using the CRT (opt variant) and baseline tokenizers, we compare results on the BLIP-3o pre-training dataset for text-to-image performance (775M auto-regressive model for 600k iterations). We show improvements on both image-text alignment and distribution matching.
Tokenizer gFID CLIPScore
Baseline 4.61 0.33
CRT (ours) 4.15 0.36
@@
Figure 13 provides a different perspective on the comparison between Figures 4 and 7. We do this by comparing stage 2 model scales directly by gFID performance by train FLOPs. Note that like the main figure, each point is an independent training run. We see that CRT tokens improve performance at all scales both for training efficiency and final performance. This difference is especially pronounced at the smaller compute and parameter scales.
Appendix G Qualitative Examples
Figure 15: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 250, class name Siberian Husky
Figure 16: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 972, class name cliff
Figure 17: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 207, class name Golden Retriever
Figure 18: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 90, class name lorikeet
Figure 19: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 29, class name axolotl
Figure 20: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 297, class name sloth bear
Figure 21: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 248, class name husky
Figure 22: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 973, class name coral reef
Figure 23: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 993, class name gyromitra
Figure 24: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 388, class name giant panda
Figure 25: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 387, class name red panda
Figure 26: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 417, class name balloon
Figure 27: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 279, class name Arctic fox
Figure 28: Non-cherrypicked generation examples from CRTopt-AR-775M (ImageNet 256x256). CFG scale = 4.0, class index 291, class name lion
@@ Appendix G Qualitative Examples
Figures 15-19 show non-cherrypicked generation examples from our best model CRTopt-AR-775M. They demonstrate a high degree of diversity and image quality across a wide array of classes. Figure 29 shows selected comparisons with and without CRT, constructed so that overall image structure is comparable. CRT demonstrates better long-range and structural consistency.
Figure 29: Selected generation comparisons with and without using the CRT tokenizer. Here we show direct comparisons between stage 2 generations with and without using the CRT tokenizer. CRT results in greater long-range structural coherence and higher quality samples. To generate these direct comparisons, we provide the first 32 (of 256) tokens from a validation example for context to the stage 2 auto-regressive model and filter for image similarity using CLIP B/32 [radford2021learning] embeddings. For more non-curated generations with our model, see Figure 14 and Appendix G.
Appendix H Text-to-image results
The main body of this work focused on ImageNet, since it is an established benchmark with good image diversity. However, much of the large-scale work on image generation focuses on text-to-image generation. To show the generality of our method, we show results in the text-to-image domain. For this experiment, we took the BLIP-3o long-caption [chen2025blip3] pre-training dataset (27M image-text pairs vs 1.2M ImageNet images), tokenized the text with OpenAI’s CLIP L/14 text tower [radford2021learning], and trained an auto-regressive image generation model conditioned on these text embeddings. This dataset is diverse, being a combination of webcrawled images and high-quality data from JourneyDB. We trained our XL sized model (775M) for 600k iterations at batch size 256 using the CRT (opt variant) and baseline tokenizers from our study. The results from this experiment are in Table 5. We use CLIPScore [hessel2021clipscore] (higher is better) with OpenAI CLIP B/32 [radford2021learning] to measure image-text alignment. We see a solid improvement with our method, even though we are re-using the image tokenizer from our ImageNet experiments and this dataset is thus OOD.
Figure 30: Reconstruction comparison (best viewed zoomed in). We show that CRT applied to baseline only minimally affects distortion, while CRT with CE introduces more noticeable blocky artifacts in reconstruction due to the strength of regularization.
Appendix I Extended Related Work
Evolution of visual tokenization. The current paradigm of discrete visual tokenizers came from VQ-VAE [van2017neural], originally introduced as a method for discrete latent representation learning in continuous domains such as image and audio. It was introduced to prevent the posterior collapse problem in VAEs by separating the compression and generation phases of training. Section 4.2 of [van2017neural] notes that their loss function leads to blurry reconstructions, and the authors suggest including perceptual metrics (metrics that align with human judgment of similarity) in the loss function. [esser2021taming] explores this idea with VQGAN, which introduces a perceptual loss based on $LPIPS$ [zhang2018unreasonable], and a GAN loss to add realistic textures. Since the advent of VQGAN, there have been numerous attempts at improving visual tokenizers [yu2021vector, gu2022rethinking, zheng2022movq, zhao2024image, yu2024image, wang2024omnitokenizer, jin2309unified, rodriguez2023ocr, wang2024image, peng2022beit, zhu2024wavelet, razzhigaev2023kandinsky]. Some works focus on improvements including ViT-VQGAN [yu2021vector], which replaces CNN layers with ViT layers, and MoVQGAN [zheng2022movq], which adds adaptive instance normalization layers in the CNN decoder. Other works focus on mitigating “codebook collapse” [huh2023straightening] in vector quantization (VQ) [gray1984vector] by introducing novel quantization schemes, for example by replacing dead codes with encoder outputs [dhariwal2020jukebox]. We refer the reader to [huh2023straightening] for a thorough review of these methods. Other works replace VQ entirely with scalar quantization techniques [mentzer2023finite], and demonstrates increased codebook utilization compared to VQ: for example, MAG-VIT [yu2023language, luo2024open] quantizes each dimension to 2 values and adds entropy loss to encourage codebook utilization.
Appendix J VQVAE comparisons
In Figure 30, we show detailed comparisons between reconstructions from baseline and various CRT configurations. CRT and baseline look quite similar, while CRT with CE (instead of L2) demonstrates some noticeable blocky artifacts as a result of the strength of regularization.